Calculate the [H3O+] of each aqueous solution with the following [OH-]:

19 of 20 Part A 0 of each aqueous solution OH stomach acid, 1.7 x 10-13 M Express your answer to two significant figures and include the appropriate units. alue Units Submit Request Answer vPart B urine, 3.8 x 109 M Express your answer to two significant figures and include the appropriate units. Value Units Submit Request Answer Part C orange juice, 2.7 x 1011 M Express your answer to two significant figures and include the appropriate units.
media%2F239%2F23907554-13ab-4fb5-a916-6b
Answer #1

We know,

Kwater = [H3O+] * [OH-]

=>  [H3O+] * [OH-] = 1*10-14

=> [H3O+] = 1*10-14 /  [OH-]

Part A:

Stomach acid - [OH-] = 1.7*10-13M

So,  [H3O+] = 1*10-14 /  [OH-] = 1*10-14 / 1.7*10-13 = 0.0588M

Therefore  [H3O+] of stomach acid = 0.059M

Part B:

Urine - [OH-] = 3.8*10-9M

So,  [H3O+] = 1*10-14 /  [OH-] = 1*10-14 /  3.8*10-9 = 2.631*10-6M

Therefore  [H3O+] of urine =2.6*10-6M

Part C:

Orange Juice - [OH-] = 2.7*10-11M

So,  [H3O+] = 1*10-14 /  [OH-] = 1*10-14 / 2.7*10-11 = 3.7*10-4M

Therefore  [H3O+] of Orange juice =3.7*10-4M

Part D:

Bile - [OH-] = 3.9*10-6M

So,  [H3O+] = 1*10-14 /  [OH-] = 1*10-14 / 3.9*10-6 = 2.56*10-9M

Therefore  [H3O+] of Bile =2.6*10-9M